OpenStudy (anonymous):

MULTIPLY: 2y(2y^2-5y+6)

OpenStudy (angela210793):

4y^3-10y^2+12y

OpenStudy (anonymous):

4y^3-10y^2+12y

OpenStudy (mathteacher1729):

To explain the answers above... We are distributing the \(2y\) (two times y) to the stuff inside the parenthesis. \[2y(2y^2-5y+6) = 2y\cdot (2y^2) - 2y\cdot (-5y)+2y\cdot (6)\] (see how 2y is multiplied by everything in the parenthesis? Now we take each expression separately. \[2y\cdot (2y^2) = 4y^3\] \[2y\cdot (-5y) = -10y^2\] \[2y\cdot(6) = 12y\] Now add everything together and you have the final answer of \[4y^3-10y^2+12y\] Hope this helps. Check out http://www.mathtv.com and look under "algebra" and then polynomials. They have lots of helpful videos there. :)